OpenStudy (anonymous):

true or false? In order to solve a system of equation by substitution, you MUST always solve for y in one of the equations and hen back-substitute

OpenStudy (anonymous):

True

OpenStudy (anonymous):

because for example if you have to unknowns and two equations, ex: y=x+5 and the other says x=y-6.....then you would have to back substitue the y=x+5 into the value for y in the second equation.

OpenStudy (anonymous):



OpenStudy (anonymous):

then just solve for x

OpenStudy (anonymous):

thank you i understand now

OpenStudy (anonymous):